# Asymptotic confidence bands for the histogram regression estimator

### v1

@@ 4 Proofs
$T_{2}=\bigg{|}\sqrt{n}\sup_{x\in[0,1]^{p}}|\tau_{\delta}(x)^{1/2}\tilde{m}^{(\varepsilon)}(x)|-\mathbf{S}_{\delta}\bigg{|}=\mathcal{O}_{\mathbb{P}}\left(\frac{(\log n)^{3/2}}{\delta^{p/2}n^{1/2-1/\nu}}+\frac{(\log n)^{5/4}}{\delta^{p/4}n^{1/4}}+\frac{\log n}{\delta^{p/6}n^{1/6}}\right)$ (22)
@@
$\frac{(\log n)^{5/2}}{\delta^{p/2}n^{1/2-1/\nu}}=\Big{(}\frac{(\log n)^{5}}{\delta^{p}n}n^{2/\nu}\Big{)}^{1/2}\to 0.$
@@
Next we consider the remainder term $T_{3}$ from (19). By the definition of $\tilde{m}^{(\varepsilon)}$ in (13) and $\hat{p}_{\delta}(x)$ in (9) we obtain
$\hat{m}^{(\varepsilon)}(x)-\tilde{m}^{(\varepsilon)}(x)=\tilde{m}^{(\varepsilon)}(x)\left(\frac{p_{\delta}(x)}{\hat{p}_{\delta}(x)}\mathbb{I}\{\hat{p}_{\delta}(x)>0\}-1\right).$
@@
$T_{3}$ $\leq$ $\left(\mathbf{S}_{\delta}+o_{\mathbb{P}}((\log n)^{-1})\right)\sup_{x\in[0,1]^{p}}\left|\frac{p_{\delta}(x)}{\hat{p}_{\delta}(x)}\mathbb{I}\{\hat{p}_{\delta}(x)>0\}-1\right|$
$=$ $\mathcal{O}_{\mathbb{P}}(\sqrt{\log n})\sup_{x\in[0,1]^{p}}\left|np_{\delta}(x)\left(\frac{1}{n\hat{p}_{\delta}(x)}-\frac{1}{np_{\delta}(x)}\right)\mathbb{I}\{\hat{p}_{\delta}(x)>0\}-\mathbb{I}\{\hat{p}_{\delta}(x)=0\}\right|.$
To obtain the remainder rate $o_{\mathbb{P}}((\log n)^{-1})$, the second indicator can be written as $(\mathbb{I}\{\exists j\in[n]:X_{j}\in A_{\delta}(x)\}-1)$ and be handled as in (20). For the other term note that for fixed $x$ the random variable $B_{n}=n\hat{p}_{\delta}(x)$ is binomially distributed with parameters $n$ and $p_{n}=p_{\delta}(x)$, and we can apply Lemma 4.1. Together with (7), the union bound and Markov’s inequality we obtain
$\mathbb{P}\left(\sup_{x\in[0,1]^{p}}\left|np_{\delta}(x)\left(\frac{1}{n\hat{p}_{\delta}(x)}-\frac{1}{np_{\delta}(x)}\right)\right|\mathbb{I}\{\hat{p}_{\delta}(x)>0\}>\frac{\kappa}{(\log n)^{3/2}}\right)$
$\leq$ $\delta^{-p}\frac{(\log n)^{3q/2}}{\kappa^{q}}(nC_{X}\delta^{p})^{q}\mathcal{O}((n\delta^{p})^{-3q/2})=\mathcal{O}\left(\delta^{-p}\frac{(\log n)^{3q/2}}{(\delta^{p}n)^{q/2}}\right)=o(1)$
@@
Altogether we have now proved
$\sqrt{n}\sup_{x\in[0,1]^{p}}\tau_{\delta}(x)^{1/2}|\hat{m}(x)-m(x)|=\mathbf{S}_{\delta}+R_{n}$ (23)
@@
$\mathbb{P}$ $\big{(}m(x)\in\mathcal{C}_{n}(x),\,\forall x\in[0,1]^{p}\big{)}$
$=\mathbb{P}\Big{(}\sqrt{n}\sup_{x\in[0,1]^{p}}\tau_{\delta}(x)^{1/2}|\hat{m}(x)-m(x)|\leq c_{\delta}(\beta)\Big{)}$
$=\mathbb{P}\big{(}\mathbf{S}_{\delta}+R_{n}\leq c_{\delta}(\beta)\big{)}$
@@
$=4\kappa_{n}\left(\mathbb{E}\left[\max_{j=1,\ldots,\delta^{-p}}|Z_{j}|\right]+1\right)$
$=\kappa_{n}\mathcal{O}(\sqrt{\log n})=o(1)$
@@
$\liminf_{n\to\infty}\inf_{m\in\mathcal{H}^{\alpha}(C_{H})}\mathbb{P}\left(m(x)\in\mathcal{C}_{n}(x),\,\forall x\in[0,1]^{p}\right)\geq 1-\beta.\qed$
@@
$\mathbb{E}\big{[}\sup_{x\in[0,1]^{p}}|\hat{m}^{(m)}(x)-m(x)|\big{]}\lesssim n^{-1/2}\delta^{-p/2}\mathbb{E}[T_{1}]\lesssim\delta^{\alpha}+\delta^{-p}(1-c_{X}\delta^{p})^{n},$
where the second term is negligible if $\log(\delta^{-1})=o(n\delta^{p})$. The latter condition is implied by assumption (4). Together with $T_{2}+T_{3}=o_{\mathbb{P}}((\log n)^{-1})$, (6) and (8), we conclude $\sup_{x\in[0,1]^{p}}|\hat{m}(x)-m(x)|=\mathcal{O}_{\mathbb{P}}(\delta^{\alpha}+\sqrt{\frac{\log n}{n\delta^{p}}})$. Choosing $\delta=(n/\log n)^{-1/(2\alpha+p)}$ balances both terms and yields the optimal rate of convergence. ∎
@@
$R_{n}$ $=$ $\bigg{|}\sqrt{n}\sup_{x\in[0,1]^{p}}\tau_{\delta}(x)^{1/2}|\hat{m}(x)-m(x)|-\sqrt{n}\sup_{x\in[0,1]^{p}}\hat{\tau}_{\delta}(x)^{1/2}|\hat{m}(x)-m(x)|\bigg{|}$
$\leq$ $\sqrt{n}\sup_{x\in[0,1]^{p}}\tau_{\delta}(x)^{1/2}|\hat{m}(x)-m(x)|\sup_{x\in[0,1]^{p}}\frac{|\hat{\tau}_{\delta}(x)^{1/2}-\tau_{\delta}(x)^{1/2}|}{\tau_{\delta}(x)^{1/2}}$
$=$ $(\mathbf{S}_{\delta}+o_{\mathbb{P}}((\log n)^{-1}))\sup_{x\in[0,1]^{p}}\frac{|\hat{\tau}_{\delta}(x)-\tau_{\delta}(x)|}{\tau_{\delta}(x)^{1/2}(\tau_{\delta}(x)^{1/2}+\hat{\tau}_{\delta}(x)^{1/2})}$
$\leq$ $\mathcal{O}_{\mathbb{P}}((\log n)^{1/2})\sup_{x\in[0,1]^{p}}\Big{|}\frac{\hat{\tau}_{\delta}(x)}{\tau_{\delta}(x)}-1\Big{|}=o_{\mathbb{P}}((\log n)^{-1})$
by (23), (25), (6) and the assumption on $\hat{\tau}_{\delta}$. ∎
@@
We will apply the result by Chernozhukov et al., 2014b (, Corollary 2.2) formulated in Theorem A.1 in the appendix to fix the notation. While the function class in this result does not depend on $n$ and $\delta$, the theorem can be applied to $\mathcal{F}_{\delta}$ for every $n$ or $\delta$, respectively, due to the universality of the occuring constants. As a consequence we get a sequence $\mathbf{S}_{\delta}$ of approximating random variables as pointed out by Chernozhukov et al., 2014b (, Remark 2.1). We consider the function class $\mathcal{F}_{\delta}$ from (15). To obtain the claimed result for the absolute value, we need to consider $\mathcal{F}_{\delta}\cup-\mathcal{F}_{\delta}$. Due to Chernozhukov et al., 2014b (, Corollary 2.1) we consider $\mathcal{F}_{\delta}$ without loss of generality.
@@
$\sup_{x\in[0,1]^{p}}|f_{x,\delta}(t,\epsilon)|$ $=\sup_{x\in[0,1]^{p}}|\epsilon|s_{\delta}(x)^{-1/2}\mathbb{I}\{t\in A_{\delta}(x)\}$
$\leq|\epsilon|(c_{X}c_{\sigma^{2}})^{-1/2}\delta^{-p/2}=:F(t,\epsilon)$ (24)
applying $c_{\sigma^{2}}c_{X}\delta^{p}\leq s_{\delta}(x)\leq C_{\sigma^{2}}C_{X}\delta^{p}$ from (10), (1) and (2). The finite size $|\mathcal{F}_{\delta}|=\delta^{-p}$ is an upper bound for any covering number of $\mathcal{F}_{\delta}$. Hence,
$\sup_{Q\in\mathcal{Q}}N(\mathcal{F}_{\delta},\|\cdot\|_{Q,2},\kappa\|F\|_{Q,2})\leq\delta^{-p}\leq\delta^{-p}/\kappa$
for all $\kappa\in(0,1]$, $\mathcal{F}_{\delta}$ is indeed of VC-type and we can choose $A=\delta^{-p}$ and $v=1$ in Theorem A.1.
@@
$\sup_{f\in\mathcal{F}_{\delta}}P|f|^{2}=\sup_{x\in[0,1]^{p}}s_{\delta}(x)^{-1}\mathbb{E}\left[\varepsilon_{1}^{2}\mathbb{I}\{X_{1}\in A_{\delta}(x)\}\right]=1,$
$\tilde{\sigma}$ from Theorem A.1 is equal to $1$. Moreover,
$\sup_{f\in\mathcal{F}_{\delta}}P|f|^{3}$ $=\sup_{x\in[0,1]^{p}}s_{\delta}(x)^{-3/2}\mathbb{E}\left[|\varepsilon_{1}\mathbb{I}\{X_{1}\in A_{\delta}(x)\}|^{3}\right]$
$\leq K(c_{\sigma^{2}}c_{X})^{-3/2}C_{X}\delta^{-p/2}.$
@@
$\|F\|_{P,\nu}=\mathbb{E}\left[|F(X_{1},\varepsilon_{1})|^{\nu}\right]^{1/\nu}=(c_{X}c_{\sigma^{2}})^{-1/2}\delta^{-p/2}\mathbb{E}\left[|\varepsilon_{1}|^{\nu}\right]^{1/\nu}.$
We thus choose $b$ from Theorem A.1 as
$b:=\delta^{-p/2}C_{b},\quad C_{b}:=\max\left\{K(c_{X}c_{\sigma^{2}})^{-3/2}C_{X},\mathbb{E}\left[|\varepsilon_{1}|^{\nu}\right]^{1/\nu}(c_{X}c_{\sigma^{2}})^{-1/2}\right\}.$
Since $b$ is of order $\delta^{-p/2}$, we have $b\geq 1$ for $\delta$ small enough. To apply Theorem A.1 we obtain
$\Gamma_{n}$ $=cv(\log n\vee\log(Ab/\tilde{\sigma}))=c(\log n\vee\log(\delta^{-p}\delta^{-p/2}C_{b}))$
$\lesssim\log n\vee p\log(\delta^{-1})\lesssim\log n.$
Let $B_{\delta}$ be the centered Gaussian process defined in the claim. For $\gamma=(\log n)^{-1}$ Theorem A.1 yields that there exists a random variable
$\mathbf{S}_{\delta}\overset{d}{=}\sup_{f\in\mathcal{F}_{\delta}\cup-\mathcal{F}_{\delta}}B_{\delta}(f)=\sup_{f\in\mathcal{F}_{\delta}}|B_{\delta}(f)|=\sup_{x\in[0,1]^{p}}|B_{\delta}(f_{x,\delta})|$
@@
$\bigg{|}$ $\sqrt{n}\sup_{x\in[0,1]^{p}}|\tau_{\delta}(x)^{1/2}\tilde{m}^{(\varepsilon)}(x)|-\mathbf{S}_{\delta}\bigg{|}$
@@
$=\mathcal{O}_{\mathbb{P}}\left(\frac{(\log n)^{3/2}}{\delta^{p/2}n^{1/2-1/\nu}}+\frac{(\log n)^{5/4}}{\delta^{p/4}n^{1/4}}+\frac{\log n}{\delta^{p/6}n^{1/6}}\right)$
since $\Gamma_{n}=\mathcal{O}(\log n)$ and $b=\mathcal{O}(\delta^{-p/2})$. Finally note, that the structure of the covariance (16) implies that
$\mathbf{S}_{\delta}\overset{d}{=}\sup_{x\in[0,1]^{p}}|B_{\delta}(f_{x,\delta})|\overset{d}{=}\max_{j=1,\ldots,\delta^{-p}}|Z_{j}|,$ (25)

### v2

@@ 6 Proofs
$T_{2}$ $=\bigg|\sqrt{n}\sup_{x\in\mathcal{X}}|\tau_{\Delta}(x)^{1/2}\widetilde{m}^{(\varepsilon)}(x)|-\mathbf{S}_{\Delta}\bigg|$
$=\mathcal{O}_{\mathbb{P}}\left(\frac{(\log n)^{3/2}\Delta^{1/2}}{n^{1/2-1/\nu}}+\frac{(\log n)^{5/4}\Delta^{1/4}}{n^{1/4}}+\frac{(\log n)\Delta^{1/6}}{n^{1/6}}\right)$ (26)
@@
$\frac{(\log n)^{5/2}\Delta^{1/2}}{n^{1/2-1/\nu}}=\Big(\frac{(\log n)^{5}\Delta}{n}n^{2/\nu}\Big)^{1/2}\to 0.$
@@
Next we consider the remainder term $T_{3}$ from (23). By the definition of $\widetilde{m}^{(\varepsilon)}$ in (17) and $\widehat{p}_{\Delta}(x)$ in (13) we obtain
$\widehat{m}^{(\varepsilon)}(x)-\widetilde{m}^{(\varepsilon)}(x)=\widetilde{m}^{(\varepsilon)}(x)\left(\frac{p_{\Delta}(x)}{\widehat{p}_{\Delta}(x)}\mathbb{I}\{\widehat{p}_{\Delta}(x)>0\}-1\right).$
@@
$T_{3}$ $\leq$ $\left(\mathbf{S}_{\Delta}+o_{\mathbb{P}}((\log n)^{-1})\right)\sup_{x\in\mathcal{X}}\left|\frac{p_{\Delta}(x)}{\widehat{p}_{\Delta}(x)}\mathbb{I}\{\widehat{p}_{\Delta}(x)>0\}-1\right|$
$=$ $\mathcal{O}_{\mathbb{P}}(\sqrt{\log n})\sup_{x\in\mathcal{X}}\left|np_{\Delta}(x)\left(\frac{1}{n\widehat{p}_{\Delta}(x)}-\frac{1}{np_{\Delta}(x)}\right)\mathbb{I}\{\widehat{p}_{\Delta}(x)>0\}-\mathbb{I}\{\widehat{p}_{\Delta}(x)=0\}\right|.$
To obtain the remainder rate $o_{\mathbb{P}}((\log n)^{-1})$, the second indicator can be written as $(\mathbb{I}\{\exists j\in[n]:X_{j}\in A_{\Delta}(x)\}-1)$ and be handled as in (24). For the other term note that for fixed $x$ the random variable $B_{n}=n\widehat{p}_{\Delta}(x)$ is binomially distributed with parameters $n$ and $p_{n}=p_{\Delta}(x)$, and we can apply Lemma 6.1. Together with (10), the union bound and Markov’s inequality we obtain
$\mathbb{P}\left(\sup_{x\in\mathcal{X}}\left|np_{\Delta}(x)\left(\frac{1}{n\widehat{p}_{\Delta}(x)}-\frac{1}{np_{\Delta}(x)}\right)\right|\mathbb{I}\{\widehat{p}_{\Delta}(x)>0\}>\frac{\kappa}{(\log n)^{3/2}}\right)$
$\leq\Delta\frac{(\log n)^{3q/2}}{\kappa^{q}}(nC_{X}C_{p}/\Delta)^{q}\mathcal{O}((n/\Delta)^{-3q/2})=\mathcal{O}\left(\Delta\frac{(\log n)^{3q/2}}{(n/\Delta)^{q/2}}\right)=o(1)$ (27)
@@
Altogether we conclude
$\sqrt{n}\sup_{x\in\mathcal{X}}\tau_{\Delta}(x)^{1/2}|\widehat{m}(x)-m(x)|=\mathbf{S}_{\Delta}+R_{n}$ (28)
@@
$\mathbb{P}$ $\big(m(x)\in\mathcal{C}_{n}(x),\,\forall x\in\mathcal{X}\big)$
$=\mathbb{P}\Big(\sqrt{n}\sup_{x\in\mathcal{X}}\tau_{\Delta}(x)^{1/2}|\widehat{m}(x)-m(x)|\leq c_{\Delta}(\beta)\Big)$
$=\mathbb{P}\big(\mathbf{S}_{\Delta}+R_{n}\leq c_{\Delta}(\beta)\big)$ (29)
@@
$=4\kappa_{n}\left(\mathbb{E}\left[\max_{j=1,\ldots,\Delta}|Z_{j}|\right]+1\right)$
$=\mathcal{O}(\kappa_{n}\sqrt{\log n})=o(1)$
@@
$\liminf_{n\to\infty}\inf_{m\in\mathcal{H}^{\alpha}(C_{H})}\mathbb{P}\left(m(x)\in\mathcal{C}_{n}(x),\,\forall x\in\mathcal{X}\right)\geq 1-\beta.\qed$
Proof of Corollary 2.4.
As in (28) and (29) we obtain
$\mathbb{P}\big(m(x)\in\widetilde{\mathcal{C}}_{n}(x),\,\forall x\in\mathcal{X}\big)$ $=$ $\mathbb{P}\big(\mathbf{S}_{\Delta}+R_{n}\leq\widetilde{c}_{\Delta}(\beta)\big)$
$=$ $\mathbb{P}\big(a_{\Delta}(\mathbf{S}_{\Delta}-b_{\Delta})+a_{\Delta}R_{n}\leq-\log(-\log(1-\beta))\big)$
$=$ $\mathbb{P}\big(a_{\Delta}(\mathbf{S}_{\Delta}-b_{\Delta})\leq-\log(-\log(1-\beta))\big)+o(1).$
The last equality follows from Slutsky’s lemma because $R_{n}=o_{\mathbb{P}}((\log n)^{-1})$, and from assumption (6) it follows that $a_{\Delta}\lesssim(\log n)^{1/2}$. It remains to verify that the limit distribution of $a_{\Delta}(\mathbf{S}_{\Delta}-b_{\Delta})\overset{d}{=}a_{\Delta}(\max_{j=1,\ldots,\Delta}|Z_{j}|-b_{\Delta})$ is the Gumbel distribution with cdf $\exp(-e^{-x})$. Indeed, using the the asymptotic independence of the maximum and minimum (Leadbetter et al.,, 2012, Theorem 1.8.2) together with the extreme value limit of the Gaussian distribution (Leadbetter et al.,, 2012, Theorem 1.5.3) we have
$\mathbb{P}(a_{\Delta}(\max_{j}|Z_{j}|-b^{\prime}_{\Delta}$ $)\leq x)=\mathbb{P}\Big(\max_{j}Z_{j}\leq\frac{x}{a_{\Delta}}+b^{\prime}_{\Delta},\min_{j}Z_{j}\geq-\frac{x}{a_{\Delta}}-b^{\prime}_{\Delta}\Big)$
$=\mathbb{P}\Big(\max_{j}Z_{j}\leq\frac{x}{a_{\Delta}}+b^{\prime}_{\Delta}\Big)\mathbb{P}\Big(\min_{j}Z_{j}\geq-\frac{x}{a_{\Delta}}-b^{\prime}_{\Delta}\Big)+o(1)$
$=\mathbb{P}\Big(a_{\Delta}(\max_{j}Z_{j}-b^{\prime}_{\Delta})\leq x\Big)\mathbb{P}\Big(-\max_{j}(-Z_{j})\geq-\frac{x}{a_{\Delta}}-b^{\prime}_{\Delta}\Big)+o(1)$
$=\mathbb{P}\Big(a_{\Delta}(\max_{j}Z_{j}-b^{\prime}_{\Delta})\leq x\Big)^{2}+o(1)\to e^{-2e^{-x}}$
for $a_{\Delta}=\sqrt{2\log\Delta}$ and $b^{\prime}_{\Delta}=\sqrt{2\log\Delta}-\frac{\log\log n+\log 4\pi}{2\sqrt{2\log\Delta}}$. Hence, $b_{\Delta}=b^{\prime}_{\Delta}+a_{\Delta}^{-1}\log 2$ yields $\mathbb{P}(a_{\Delta}(\max|Z_{j}|-b_{\Delta})\leq x)\to e^{-e^{-x}}$. ∎
@@
$\mathbb{E}\big[\sup_{x\in\mathcal{X}}|\widehat{m}^{(m)}(x)-m(x)|\big]\lesssim n^{-1/2}\Delta^{1/2}\mathbb{E}[T_{1}]\lesssim\Delta^{-\alpha/p}+\Delta(1-c_{X}c_{p}/\Delta)^{n},$
where the second term is negligible if $\log(\Delta)=o(n/\Delta)$. The latter condition is implied by assumption (6). Together with $T_{2}+T_{3}=o_{\mathbb{P}}((\log n)^{-1})$, (9) and (11), we conclude $\sup_{x\in\mathcal{X}}|\widehat{m}(x)-m(x)|=\mathcal{O}_{\mathbb{P}}(\Delta^{-\alpha/p}+\sqrt{(\log n)\Delta/n})$. Choosing $\Delta=(n/\log n)^{1/(2\alpha/p+1)}$ balances both terms and yields the optimal rate of convergence. ∎
@@
$R_{n}$ $=\bigg|\sqrt{n}\sup_{x\in\mathcal{X}}\tau_{\Delta}(x)^{1/2}|\widehat{m}(x)-m(x)|-\sqrt{n}\sup_{x\in\mathcal{X}}\widehat{\tau}_{\Delta}(x)^{1/2}|\widehat{m}(x)-m(x)|\bigg|$
$\leq\sqrt{n}\sup_{x\in\mathcal{X}}\tau_{\Delta}(x)^{1/2}|\widehat{m}(x)-m(x)|\sup_{x\in\mathcal{X}}\frac{|\widehat{\tau}_{\Delta}(x)^{1/2}-\tau_{\Delta}(x)^{1/2}|}{\tau_{\Delta}(x)^{1/2}}$
$=\big(\mathbf{S}_{\Delta}+o_{\mathbb{P}}((\log n)^{-1})\big)\sup_{x\in\mathcal{X}}\frac{|\widehat{\tau}_{\Delta}(x)-\tau_{\Delta}(x)|}{\tau_{\Delta}(x)^{1/2}(\tau_{\Delta}(x)^{1/2}+\widehat{\tau}_{\Delta}(x)^{1/2})}$
$\leq\mathcal{O}_{\mathbb{P}}((\log n)^{1/2})\sup_{x\in\mathcal{X}}\Big|\frac{\widehat{\tau}_{\Delta}(x)}{\tau_{\Delta}(x)}-1\Big|=o_{\mathbb{P}}((\log n)^{-1})$
by (28), (32), (9) and the assumption on $\widehat{\tau}_{\Delta}$. ∎
@@
We will apply the result by Chernozhukov et al., 2014b (, Corollary 2.2) formulated in Theorem B.1 in the appendix to fix the notation. While the function class in this result does not depend on $n$ and $\Delta$, the theorem can be applied to $\mathcal{F}_{\Delta}$ for every $n$ or $\Delta$, respectively, due to the universality of the occuring constants. As a consequence we get a sequence $\mathbf{S}_{\Delta}$ of approximating random variables as pointed out by Chernozhukov et al., 2014b (, Remark 2.1). We consider the function class $\mathcal{F}_{\Delta}$ from (19). To obtain the claimed result for the absolute value, we need to consider $\mathcal{F}_{\Delta}\cup-\mathcal{F}_{\Delta}$. Due to Chernozhukov et al., 2014b (, Corollary 2.1) we consider $\mathcal{F}_{\Delta}$ without loss of generality.
@@
$\sup_{x\in\mathcal{X}}|f_{x,\Delta}(t,\epsilon)|$ $=\sup_{x\in\mathcal{X}}|\epsilon|s_{\Delta}(x)^{-1/2}\mathbb{I}\{t\in A_{\Delta}(x)\}$
$\leq|\epsilon|(c_{X}c_{\sigma^{2}}c_{p})^{-1/2}\Delta^{1/2}=:F(t,\epsilon)$ (30)
applying $c_{\sigma^{2}}c_{X}c_{p}\Delta^{-1}\leq s_{\Delta}(x)\leq C_{\sigma^{2}}C_{X}C_{p}\Delta^{-1}$ from (14), (1), (2) and Assumption 2.1. The finite size $|\mathcal{F}_{\Delta}|=\Delta$ is an upper bound for any covering number of $\mathcal{F}_{\Delta}$. Hence,
$\sup_{Q\in\mathcal{Q}}N(\mathcal{F}_{\Delta},\|\cdot\|_{Q,2},\kappa\|F\|_{Q,2})\leq\Delta\leq\Delta/\kappa$ (31)
for all $\kappa\in(0,1]$, $\mathcal{F}_{\Delta}$ is indeed of VC-type and we can choose $A=\Delta$ and $v=1$ in Theorem B.1.
@@
$\sup_{f\in\mathcal{F}_{\Delta}}P|f|^{2}=\sup_{x\in\mathcal{X}}s_{\Delta}(x)^{-1}\mathbb{E}\left[\varepsilon_{1}^{2}\mathbb{I}\{X_{1}\in A_{\Delta}(x)\}\right]=1,$
the quantity $\widetilde{\sigma}$ from Theorem B.1 is equal to $1$. Moreover,
$\sup_{f\in\mathcal{F}_{\Delta}}P|f|^{3}$ $=\sup_{x\in\mathcal{X}}s_{\Delta}(x)^{-3/2}\mathbb{E}\left[|\varepsilon_{1}\mathbb{I}\{X_{1}\in A_{\Delta}(x)\}|^{3}\right]$
$\leq K(c_{\sigma^{2}}c_{X}c_{p})^{-3/2}C_{X}C_{p}\Delta^{1/2}.$
@@
$\|F\|_{P,\nu}=\mathbb{E}\left[|F(X_{1},\varepsilon_{1})|^{\nu}\right]^{1/\nu}=(c_{X}c_{\sigma^{2}}c_{p})^{-1/2}\Delta^{1/2}\mathbb{E}\left[|\varepsilon_{1}|^{\nu}\right]^{1/\nu}.$
We thus choose $b$ from Theorem B.1 as
$b:=\Delta^{1/2}C_{b},\quad C_{b}:=\max\left\{K(c_{X}c_{\sigma^{2}}C_{p})^{-3/2}C_{X}C_{p},\mathbb{E}\left[|\varepsilon_{1}|^{\nu}\right]^{1/\nu}(c_{X}c_{\sigma^{2}}C_{p})^{-1/2}\right\}.$
Since $b$ is of order $\Delta^{1/2}$, we have $b\geq 1$ for $\Delta$ large enough. To apply Theorem B.1 we obtain
$\Gamma_{n}$ $=cv(\log n\vee\log(Ab/\widetilde{\sigma}))=c(\log n\vee\log(\Delta^{3/2}C_{b}))$
$\lesssim\log n\vee p\log(\Delta)\lesssim\log n.$
Let $B_{\Delta}$ be the centered Gaussian process defined in the claim. For $\gamma=(\log n)^{-1}$ Theorem B.1 yields that there exists a random variable
$\mathbf{S}_{\Delta}\overset{d}{=}\sup_{f\in\mathcal{F}_{\Delta}\cup-\mathcal{F}_{\Delta}}B_{\Delta}(f)=\sup_{f\in\mathcal{F}_{\Delta}}|B_{\Delta}(f)|=\sup_{x\in\mathcal{X}}|B_{\Delta}(f_{x,\Delta})|$
@@
$\bigg|$ $\sqrt{n}\sup_{x\in\mathcal{X}}|\tau_{\Delta}(x)^{1/2}\widetilde{m}^{(\varepsilon)}(x)|-\mathbf{S}_{\Delta}\bigg|$
@@
$=\mathcal{O}_{\mathbb{P}}\left(\frac{(\log n)^{3/2}\Delta^{1/2}}{n^{1/2-1/\nu}}+\frac{(\log n)^{5/4}\color[rgb]{0,0,1}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,1}\Delta^{1/4}}{n^{1/4}}+\frac{(\log n)\Delta^{1/6}}{n^{1/6}}\right)$
since $\Gamma_{n}\lesssim\log n$ and $b\lesssim\Delta^{1/2}$. Finally note, that the structure of the covariance (20) implies that
$\mathbf{S}_{\Delta}\overset{d}{=}\sup_{x\in\mathcal{X}}|B_{\Delta}(f_{x,\Delta})|\overset{d}{=}\max_{j=1,\ldots,\Delta}|Z_{j}|,$ (32)